please help me solve this question:
clatheria is 40 years elder than clare. 5 years ago clatheria was 3 times as old as clare. what is their age now?

Found 2 solutions by Cromlix, fractalier:
Answer by Cromlix(4381)

Hi there,
clatheria is 40 years elder than clare
(Ct = Clatheria: Cl = Claire)
Ct = 40 + Cl........(1)
5 years ago clatheria was 3 times as old as clare.
Ct - 5 = 3(Cl - 5)......(2)
Replace Ct in Eq(2) with
(40 + Cl) from Eq(1)
40 + Cl - 5 = 3(Cl - 5)
40 + Cl - 5 = 3Cl - 15
Collect like terms:
Cl - 3Cl = -15 + 5 - 40
-2Cl = -50
2Cl = 50
Cl = 25
Claire is 25 years old
Clatheria is 65 years old.
Proof:
5 years ago clatheria was 3 times as old as clare.
65 - 5 = 3(25 - 5)
60 = 3(20)
Hope this helps :-)

Answer by fractalier(6550)

Call Clatheria's and Clare's ages NOW, x and y, respectively.
Then we write
x = y + 40
Five years ago
x - 5 = 3(y - 5)
Now substitute and get
y+40 - 5 = 3(y - 5)
y + 35 = 3y - 15
Add 15 and subtract y and get
y + 50 = 3y
2y = 50
y = 25 (Clare) and then
x = y + 40 = 65 (Clatheira)